OpenStudy (anonymous):

How do I solve for x? -4(x+10)-6=-3(x-2)

OpenStudy (anonymous):

First expand. The -4 distributes to the x and the 10 on the left and the -3 distributes to the x and the -2 on the right.

OpenStudy (anonymous):

Then what??

OpenStudy (anonymous):

Then you collect like terms, move all the x's to one side and all the numbers to the other. What do you get after expanding?

OpenStudy (anonymous):

-4x+-40-6=-3x+6

OpenStudy (anonymous):

Very good! Now the -40 and the -6 combine on the left and you get -4x-46=-3x+6 Now you want to collect like terms. So move all the x's to the left and move all the numbers to the right. I'll do the first one. To move the x's to the left, I will add 3x on both sides -4x-46=-3x+6 +3x +3x -x -46=6